OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

isnt that one of those y=mx+b questions

OpenStudy (anonymous):

It's my last question for segment one. D;

OpenStudy (anonymous):

oh lucky i need to finish segment one :/

OpenStudy (anonymous):

Yes, You'll have to first make all the answer choices look like "y=mx+b" I could do it but it will take about 10 minutes

OpenStudy (anonymous):

solve this system: -2x + h = 1 x + h = 10 after you solve it, you will get the equation of line in the form: y = mx + h

OpenStudy (anonymous):

-3x + 2h = 11?

OpenStudy (anonymous):

its not that hard all you have to do is find out whats m x1 and y1 and write out your problem

OpenStudy (anonymous):

@Keridavis can you attempt it?

OpenStudy (anonymous):

I can yes ma'am

OpenStudy (anonymous):

let's take the second equation: x = 10 - h substitute on the first one: -2(10-h)+h = 1 -20 + 2h + h = 1 3h = 21 h = 7 Now plug it back on second equation: x = 10 - 7 x = 3 so : y = 3x + 7

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3x+-+y+%3D+-7

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3x+-+y+%3D+ -5

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3x+-+y+%3D+5

OpenStudy (anonymous):

3x - y = -5?

OpenStudy (anonymous):

or 3x - y = 5?

OpenStudy (anonymous):

3x - y = -7

OpenStudy (anonymous):

A. Y = 3x +7 B. Y = 3x +5 C. Y = 3x -5 D. 1/3 (-x-5)

OpenStudy (anonymous):

3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

@M4thM1nd are you sure thats the correct answer?

OpenStudy (anonymous):

i agree with M4thM1nd

OpenStudy (anonymous):

100% sure

OpenStudy (anonymous):

you can always plug this two given points on that equation and you will see that, indeed it's the right answer